OpenStudy (anonymous):

Terry is mailing two boxes. Together they weigh 21 lbs. If the smaller box is 5 lbs. less than the larger one, how much does each box weigh? Write and solve an equation that models this scenario

OpenStudy (marziajackson$):

are there any choices?

OpenStudy (marziajackson$):

the answer is 8

OpenStudy (anonymous):

no thats all the question says

OpenStudy (anonymous):

i dont wanna be greedy but can u write th equation out, im horrible at math.lol

OpenStudy (marziajackson$):

sure :)

OpenStudy (marziajackson$):

x + (x-5) =21, (x-5) representing the smaller one. 2x-5=21 2x=26 x=13 (x-5)=13 x=8

OpenStudy (anonymous):

ur awesome

OpenStudy (anonymous):

wanna do more? ur like a genius

OpenStudy (marziajackson$):

sure and thank you

OpenStudy (anonymous):

The length of a pool is 3 feet more than twice its width. If the perimeter of the pool is 72 feet, find the dimensions of the pool by writing and solving a system of equations.

OpenStudy (anonymous):

ur 14 and this smart? what are they teaching u kids lol

OpenStudy (marziajackson$):

lol

OpenStudy (marziajackson$):

Width = W Ft. Length = (2W + 3)Ft P = 2W + 2(2W+3) = 72 2W + 4W+6 = 72 6W = 72-6 = 66 W = 11 Ft. Length = 2*11 + 3 = 25 Ft

OpenStudy (anonymous):

geez!! what grade are u in

OpenStudy (anonymous):

A desk and a chair cost $200 as a set. If the desk costs four times more than the chair, how much does each one cost? Write and solve using a system of equations.

OpenStudy (marziajackson$):

7th grade

OpenStudy (anonymous):

what the heck! now i feel dumb

OpenStudy (marziajackson$):

d = desk c = chair Equation: d + c = 200 d = 4c

OpenStudy (anonymous):

Rachel has 15 coins with a value of $2.85. If the coins are either dimes or quarters, how many of each coin does she have? Write and solve using a system of equations

OpenStudy (anonymous):

thanks again